OpenStudy (anonymous):

It is possible to project a particle with a given velocity in 2 possible ways so as to make it pass through a pt P at a distance r from the point of projection. The product of the times taken to reach this point in two possible ways is then proportional to Options 1/r r r^3 1/(r)^2

OpenStudy (anonymous):

range=u^2sin2@/g time=2usin@/g for complimentary angles we get the same range like 30 and 60 so sin30=cos60 use these infos to proceed i will help if u get stuck

OpenStudy (anonymous):

thanks a lot now is this correct? t*T=2u^2sin2@/g^2 so t*T/r=2/g and t*T is proprtional to r??